All firms have the same MC curve; it always costs $5 to make each jar of honey. Demand for honey is D:P= 20-0.05Q

What are the market equilibrium and quantity?
1
Expert's answer
2017-05-08T08:55:09-0400
TR=P*Q=20Q-0.05Q^2

MR=20-0.1Q

20-0.1Q=5

Q=150

P=12.5